A merchant has coffee worth $20 a pound that she wishes to mix with 50 pounds of coffee worth $80 a pound to get a mixture that can be sold for $60 a pound. How many pounds of the $20 coffee should be used?
Answer by mananth(16949)

A merchant has coffee worth $20 a pound that she wishes to mix with 50 pounds of coffee worth $80 a pound to get a mixture that can be sold for $60 a pound. How many pounds of the $20 coffee should be used?
$20 coffee---------------------- $80 coffee------------------------- mix $60
x-50 ---------------------------------50------------------------------- x
(x-50)*20 +50*80 =60x
20x-1000 +4000 = 60x
20x-60x= -4000 +1000
-40x= -3000
x= -3000/-40
75 The mix will be 75 lbs
And $20 coffee will be 75 -50 = 25lbs